OpenStudy (anonymous):

how many solutions exist for the system x + y = 1 and y = -x - 2

OpenStudy (anonymous):

rewrite the left one as y = 1 - x and then you have y = 1-x and y = -x -2 solutions occur when these two lines intersect, since they are straight lines they will intersect only once